# Impact of spatial curvature on quantum Otto engines

### v1

@@ Abstract
In this paper, we consider a quantum Otto cycle with a quantum harmonic oscillator on a circle as its working substance. Since the eigen-energies of this oscillator depend on the curvature of the circle, this model, as an analog model, enables us to investigate the curvature effects of the physical space on properties of quantum heat engines. We assume that two classical hot and cold thermal baths are located at places with different curvatures. We calculate the curvature-dependent work and heat in our Otto cycle with a particular emphasis on how curvature affects it’s thermal efficiency. By changing the curvature difference between the bath’s locations, we find that the efficiency of our heat engine can exceed the Carnot efficiency limit.
@@
Spatial curvature, quantum harmonic oscillator on a circle, quantum Otto cycle.
I Introduction
@@
Einstein’s theory of general relativity indicates that massive objects change the geometric nature, and thus, the curvature of the surrounding space-time, leading to some curvature-dependent effects [1, 2]. Furthermore, the equivalence principle provides a clear insight into the impact of classical gravitational fields on specific physical experiments [3]. Although these curvature effects predicted by general relativity are tremendous on cosmological scales, their influences in a laboratory on the earth are weak and difficult to detect. This is where the concept of analog models becomes relevant [4, 5, 6, 7]. Analog models have gained prominence in physics and mathematics, offering new perspectives and enabling the exchange of ideas between different scientific disciplines [8]. In recent years, there have been numerous attempts to employ physical systems, as analogy platforms to simulate and explore some aspects of general relativity theory in laboratory [9, 10, 11, 12]. For instance, analog models have been used to understand black hole physics, leading to the development of a thermodynamic theory for these objects and the discovery of the Hawking effect [13]. Furthermore, several types of analog physical systems on two-dimensional curved surfaces have been proposed to model the general relativity. Given a constant time and extracting the equatorial slice of Friedmann-Robertson-walker’s space-time, we arrive at a two-dimensional curved surface with constant curvature. In these types of analog models, a curved space is created by engineering the geometry of the space itself to investigate its effects on certain physical systems [14, 15].
On the other hand, quantum thermodynamics [16, 17] has attracted considerable attention for elucidating the fundamental relationship between quantum mechanics and thermodynamics [18]. The conversion of energy into mechanical work is essential for almost any industrial process [19]. Quantum thermodynamics investigates fundamental concepts such as temperature, heat, and work in the context of quantum physics. This field primarily focuses on studying thermal machines that operate in the quantum realm and determining how quantum properties can be exploited to enhance their efficiency. Quantum heat engines, which attracted much attention in recent year, are designed based on quantum thermodynamic cycles with different working substances, such as spin systems [20], single ions [21], harmonic-oscillators [22, 23, 24], quantum Brownian oscillators [25], two-level systems[26, 27, 28], single-particles [29, 30, 31], two-particles [32], multilevel systems [33, 34], optomechanical systems [35], and cavity quantum electrodynamics systems [36]. Among these, the so-called quantum Otto engine is presented which consists of a quantum working substance manipulated between two heat reservoirs under two adiabatic and two isochoric branches [37, 38, 39]
@@
To address the question of the efficiency of a quantum heat engine, compared to its classical counterpart, several protocols have been proposed and analyzed[40]. In [41], it is shown that non-thermal baths, such as coherent bath [42] and squeezed bath [43], may impart not only heat, but also mechanical work to a heat engine. While such engines can exceed the Carnot efficiency, they do not violate the second-law of thermodynamics. Although the Carnot bound is inapplicable to such a hybrid machine, recent proposals suggest that a small scale heat engine may potentially exceed Carnot efficiency without requiring additional resources like coherence or correlations [44, 45].
@@
In Ref.[46], by using a delta-coupled Unruh-Dewitt detector in Minkowski space-time, a relativistic quantum Otto engine is considered to extract work from a quantum field in a thermal state. The performance of a quantum Otto engine with a working medium of a relativistic particle is analyzed in Ref.[47]. Additionally, the thermal properties of a quantum heat engine with a relativistic moving heat bath are investigated in Ref.[48]. In a related study, a quantum Otto engine was introduced, in which the effective temperatures observed by stationary moving observers are utilized [49]. Also, the amount of work extracted by an observer moving in a circular path is investigated in Ref.[50]. Further research has investigated how relativistic energies and space-time geometry affect the thermal efficiency of thermodynamic cycles [51]. However, the impact of curvature on the performance of quantum thermal engines remains a relatively unexplored topic.
@@
In the present contribution, we adopt an analog model of general relativity to investigate the effects of spatial curvature on a quantum Otto engine. We consider a quantum Otto cycle with a quantum harmonic oscillator on a circle as it’s working substance. We assume that two classical hot and cold thermal baths are located at places with different curvatures. We calculate the curvature-dependent work and heat in our Otto cycle and study the curvature effects on it’s thermal efficiency. It is worth noting that the properties of the harmonic oscillator on a circle depend on the curvature of the circle, and therefore, making it a suitable platform to investigate the spatial curvature effects on this physical process. Moreover, this analog model has the advantage of being analytically tractable.
@@
The paper is organized as follows. In Section II, we briefly review the quantum harmonic oscillator on a circle. In Section III, after a short review of the quantum Otto’s engines, we propose our curvature-dependent scheme of a quantum Otto engine with a harmonic oscillator on a circle as the working substance. Section IV is devoted to investigating the efficiency of our engine, as well as its limiting case of small curvature (near the Earth surface) and large curvature (near massive celestial bodies like black holes). Then, equipped with the theoretical tools presented in previous sections, we present our results about the curvature effects on the efficiency of our quantum Otto’s engine in Section V. Finally the summary and concluding remarks are given in Section VI.
II Quantum Harmonic Oscillator on Circle
@@
Recently, some of the authors of this paper have studied the quantum dynamics of a harmonic oscillator constrained on a circle with radius $R$ [52]. By using the gnomonic projection, which is a projection onto the tangent line from the center of the circle and denoting the Cartesian coordinate of this projection by $x$, we obtain the Hamiltonian of the quantum harmonic oscillator on a circle as follows:
$\hat{H}({\lambda})=\frac{1}{2}[-(1+\lambda x^{2})^{2}\frac{d^{2}}{dx^{2}}-2\lambda x(1+\lambda x^{2})\frac{d}{dx}]+\frac{1}{2}x^{2},$ (1)
@@
It is worth noting that the spatial curvature parameter, which introduces a nonlinear character to the quantum oscillator on a circle, is reflected in the anharmonicity of the energy spectrum. In other words, when we go from a flat line to a curved circle, the energy levels are no longer evenly spaced. Therefore, our quantum oscillator exhibits an effective piston-like behavior through variations of the curvature parameter $\lambda$.
In Fig. 2, we have plotted the dimensionless variation of energy gaps:
@@
as a function of the spatial curvature parameter. It is seen that for a fixed value of $\lambda$ $(\lambda\neq 0)$, the energy gaps increase with increasing the value of $n$. Furthermore, with increasing the curvature $\lambda$, the uniformity of the energy gaps between the eigen-energies is further lost.
III QUANTUM CURVATURE-DEPENDENT OTTO ENGINE
In this section, we consider a quantum Otto’s engine with a harmonic oscillator on a circle as its working substance, as illustrated in Fig. 3. This engine consists of two isochoric branches, one with a hot bath and the other with a cold thermal bath at two different temperatures $T_{h}>T_{c}$ [53], [54]. A quantum isochoric process (constant volume process) is similar to its classical isochoric process. In contrast, a classical adiabatic thermodynamic process does not always require the occupation probabilities to remain constant, while a quantum adiabatic process proceeds slowly enough so that the generic quantum adiabatic condition is satisfied, and keeps the population distributions constant [55]. Moreover, during quantum adiabatic stages, the working substance Hamiltonian varies while population distribution remains constant [56]. It is worth mentioning that if this variation in the Hamiltonian causes the energy gaps to scale at a constant ratio, the efficiency of the engine is like its classical counterpart. However, if the deformation of the Hamiltonian leads to inhomogeneous energy gaps, the heat engine’s efficiency can exceed its classical counterpart [57].
@@ III QUANTUM CURVATURE-DEPENDENT OTTO ENGINE
The first step in understanding the thermodynamics of a harmonic oscillator on a circle is to find the correct expression for the equilibrium distribution when the system is in contact with a thermal bath. The equilibrium distribution of a quantum oscillator on a circle, weakly coupled to a large classical bath, will eventually reach the corresponding Gibbs thermal equilibrium state, $\rho=\frac{e^{-\hat{H}(\lambda)/T}}{Z(\lambda)}$, [58] where the canonical partition function of the quantum oscillator, acting as the working substance of our Otto’s engine, is given by
@@
III.1 Processes of a quantum Otto engine cycle
@@ 1. Hot isochore $(A\rightarrow B)$:
$Q_{h}=\sum_{n}E_{n}(\lambda)dP_{n}(\lambda)=\sum_{n}E_{n}^{A}(\lambda_{2})[P_{n}^{B}(\lambda_{2})-P_{n}^{A}(\lambda_{2})].$ (8)
@@
2. Isentropic compression $(B\rightarrow C)$:
@@
$W_{B\rightarrow C}=\sum_{n}P_{n}dE_{n}=\sum_{n}P_{n}^{B}(\lambda_{2})[E_{n}^{A}(\lambda_{2})-E_{n}^{C}(\lambda_{1})].$ (9)
3. Cold isochore $(C\rightarrow D)$:
@@
$Q_{c}=\sum_{n}E_{n}^{C}(\lambda_{1})[P_{n}^{D}(\lambda_{1})-P_{n}^{C}(\lambda_{1})],$ (10)
@@
4. Isentropic expansion $(D\rightarrow A)$:
@@
$W_{D\rightarrow A}=\sum_{n}P_{n}^{D}(\lambda_{2})[E_{n}^{D}(\lambda_{2})-E_{n}^{A}(\lambda_{1})].$ (11)
@@
III.2 Work and heat of the quantum Otto cycle
Now we can obtain the net work performed by our quantum Otto engine [60] as follows:
$W$ $=$ $-(W_{C\rightarrow B}+W_{D\rightarrow A})=Q_{h}-Q_{c}$
$=$ $\sum_{n}[E_{n}^{A}(\lambda_{2})-E_{n}^{C}(\lambda_{1})][P_{n}^{B}(\lambda_{2})-P_{n}^{A}(\lambda_{2})].$
@@ III.2 Work and heat of the quantum Otto cycle
IV THE THERMAL EFFICIENCY OF QUANTUM OTTO ENGINE
The first law of thermodynamics explains how to convert heat, which is a form of energy, into work. The second law tells how much of this heat can be converted into work. Efficiency is defined as the ratio of output energy to input energy [61]. The thermal efficiency of our Otto cycle is defined by
@@ IV THE THERMAL EFFICIENCY OF QUANTUM OTTO ENGINE
IV.1 LIMITING CASE OF THERMAL EFFICIENCY
@@
1. The small spatial curvature limit:
@@
2. The large spatial curvature limit:
@@
$Q_{h}=\frac{\lambda}{2}[\frac{q_{h}\theta^{\prime}_{3}(q_{h})}{\theta_{3}(q_{h})-1}-\frac{q_{c}\theta^{\prime}_{3}(q_{c})}{\theta_{3}(q_{c})-1}],$
@@
V RESULTS AND REMARKS
In this section, we proceed to study the effects of curvature on our quantum Otto’s engine using the analog model of a simple quantum harmonic oscillator on a circle. As shown in Eq. (III.2), the work extracted and thus the heat engine’s efficiency of our proposed quantum Otto’s heat engine depend on the net variation of the energy eigenvalues between points $A$ and $C$, as well as the population difference between points $A$ and $B$.
@@ V RESULTS AND REMARKS
One of the important factors in enhancing the efficiency of a quantum heat engine over its classical counterpart is that the change in the Hamiltonian of it’s working substance should be such that the difference between the energy gaps during the engine’s irreversible process changes in an anharmonicity manner [57].
It is worth noting that the results of Fig. 9 indicate that the efficiency of our analog heat engine model depends on the spatial curvature of the working substance’s location. Therefore, by utilizing a quantum harmonic oscillator on a circle as the working substance, we can control the work and efficiency of the Otto engine. Moreover, it is possible to enhance it’s efficiency to an upper limit that is greater than the Carnot efficiency bound $(\eta_{Carnot}=1-\frac{T_{cold}}{T_{hot}})$, by adjusting the curvature of space parameters $\lambda_{1}$ and $\lambda_{2}$ (the height of the hot and cold baths). The inset of Fig. 9 clearly shows that the efficiency of our Otto engine exceeds the Carnot efficiency limit.
Finally, to better visualize the results that have been obtained, we display in Fig. 10 the variation of our engine efficiency versus the curvatures $\lambda_{1}$ and $\lambda_{2}$
VI SUMMARY AND CONCLUDING REMARKS
In this paper, we have considered a quantum Otto’s engine with a quantum harmonic oscillator on a circle as it’s working substance, serving as an analog model of general relativity. By utilizing the curvature-dependent energies of this oscillator, we have investigated how the curvature effects of the physical space affects the properties of our quantum heat engine. Assuming that two thermal baths are located at places with different curvatures, we have calculated the curvature-dependent work and heat, with a particular focus on the effects of curvature on it’s thermal efficiency. The result show that by adjusting the curvature difference between the bath’s locations, we can control the efficiency of our heat engine, potentially allowing it to exceed the Carnot efficiency limit.
Acknowledgements.

### v2

@@ Abstract
In this paper, we consider a quantum Otto cycle with a quantum harmonic oscillator on a circle as its working substance. Since the eigen-energies of this oscillator depend on the curvature of the circle, this model, as an analog model, enables us to investigate the curvature effects of the physical space on properties of quantum heat engines. We assume that two classical hot and cold thermal baths are located at places with different curvatures. We calculate the curvature-dependent work and heat in our Otto cycle with a particular emphasis on how curvature affects it’s thermal efficiency. By adjusting the curvature difference between the locations of the thermal baths, we demonstrate that the efficiency of our heat engine can reach the Carnot limit.
keywords:
@@
Spatial curvature, quantum harmonic oscillator on a circle, quantum Otto cycle
1 Introduction
@@
Einstein’s theory of general relativity indicates that massive objects change the geometric nature, and thus, the curvature of the surrounding space-time, leading to some curvature-dependent effects [1, 2]. Furthermore, the equivalence principle provides a clear insight into the impact of classical gravitational fields on specific physical experiments [3]. Although these curvature effects predicted by general relativity are tremendous on cosmological scales, their influences in a laboratory on the earth are weak and difficult to detect. This is where the concept of analog models becomes relevant [4, 5, 6, 7]. Analog models have gained prominence in physics and mathematics, offering new perspectives and enabling the exchange of ideas between different scientific disciplines [8]. In recent years, there have been numerous attempts to employ physical systems, as analogy platforms to simulate and explore some aspects of general relativity theory in laboratory [9, 1, 10, 11, 12]. For instance, analog models have been used to understand black hole physics, leading to the development of a thermodynamic theory for these objects and the discovery of the Hawking effect [13]. Furthermore, several types of analog physical systems on two-dimensional curved surfaces have been proposed to model the general relativity. Given a constant time and extracting the equatorial slice of Friedmann-Robertson-walker’s space-time, we arrive at a two-dimensional curved surface with constant curvature. In these types of analog models, a curved space is created by engineering the geometry of the space itself to investigate its effects on certain physical systems [16, 14, 15]. On the other hand, quantum thermodynamics [17, 18] has attracted considerable attention for elucidating the fundamental relationship between quantum mechanics and thermodynamics [19]. The conversion of energy into mechanical work is essential for almost any industrial process [20]. Quantum thermodynamics investigates fundamental concepts such as temperature, heat, and work in the context of quantum physics. This field primarily focuses on studying thermal machines that operate in the quantum realm and determining how quantum properties can be exploited to enhance their efficiency. Quantum heat engines, which attracted much attention in recent year, are designed based on quantum thermodynamic cycles with different working substances, such as spin systems [21, 22], single ions [19], harmonic-oscillators [23, 24, 25], quantum Brownian oscillators [26], two-level systems [27, 28, 29], single-particles [30, 31, 32], two-particles [33], multilevel systems [34, 35], optomechanical systems [36], and cavity quantum electrodynamics systems [37]. Among these, the so-called quantum Otto engine is presented, consisting of a quantum working substance manipulated between two heat reservoirs under two adiabatic and two isochoric branches [38, 39, 40].
@@
To address the question of the efficiency of a quantum heat engine, compared to its classical counterpart, several protocols have been proposed and analyzed [41]. In [42], it is shown that non-thermal baths, such as coherent bath [43] and squeezed bath [44], may impart not only heat, but also mechanical work to a heat engine. While such engines can exceed the Carnot efficiency, they do not violate the second-law of thermodynamics[45].
@@
In Ref. [15], by using a delta-coupled Unruh-Dewitt detector in Minkowski space-time, a relativistic quantum Otto engine is considered to extract work from a quantum field in a thermal state. In Ref. [46], quantum Otto refrigeration cycles are constructed by utilizing the gravitational redshift experienced by photons in curved spacetime. The performance of a quantum Otto engine with a working medium of a relativistic particle is analyzed in Ref. [47]. Additionally, the thermal properties of a quantum heat engine with a relativistic moving heat bath are investigated in Ref. [48]. In a related study, a quantum Otto engine was introduced, in which the effective temperatures observed by stationary moving observers are utilized [49]. Also, the amount of work extracted by an observer moving in a circular path is investigated in Ref. [50]. Further research has investigated how relativistic energies and space-time geometry affect the thermal efficiency of thermodynamic cycles [51]. However, the impact of curvature on the performance of quantum thermal engines remains a relatively unexplored topic.
@@
In the present contribution, we adopt an analog model of general relativity to investigate the effects of spatial curvature on a quantum Otto engine. We consider a quantum Otto cycle with a quantum harmonic oscillator on a circle as it’s working substance. We assume that two classical hot and cold thermal baths are located at places with different curvatures. We calculate the curvature-dependent work and heat in our Otto cycle and study the curvature effects on it’s thermal efficiency. It is worth noting that the properties of the harmonic oscillator on a circle depend on the curvature of the circle, and therefore, making it a suitable platform to investigate the spatial curvature effects on this thermodynamic process. Moreover, this analog model offers the advantage of analytical tractability, facilitating a detailed exploration of curvature-induced phenomena.
@@
The paper is organized as follows. In Section II, we provide a brief review of the quantum harmonic oscillator on a circle. In Section III, after a concise overview of the quantum Otto’s engines, we propose our curvature-dependent scheme of a quantum Otto engine with a harmonic oscillator on a circle as the working substance. Section IV is devoted to investigating the efficiency of our engine, as well as its limiting case of small curvature (near the Earth surface) and large curvature (near massive celestial bodies like black holes). In Section V, utilizing the theoretical framework established in the previous sections, we present our findings on the impact of curvature on the efficiency of our quantum Otto’s engine. Finally the summary and concluding remarks are given in Section VI.
2 Quantum Harmonic Oscillator on Circle
@@
Recently, some of the authors of this paper studied the quantum dynamics of a harmonic oscillator constrained on a circle with radius $R$ [52]. By employing the gnomonic projection, which is a projection onto the tangent line from the center of the circle and denoting the Cartesian coordinate of this projection by $x$ (see Fig. 1), the Hamiltonian of the quantum harmonic oscillator on a circle is derived as follows:
$\hat{H}({\lambda})=\frac{1}{2}\Big{[}-(1+\lambda x^{2})^{2}\frac{d^{2}}{dx^{2}}-2\lambda x(1+\lambda x^{2})\frac{d}{dx}\Big{]}+\frac{1}{2}x^{2},$ (1)
@@
It is worth noting that the spatial curvature parameter, which introduces a nonlinear character to the quantum oscillator on a circle, manifests in the anharmonicity of the energy spectrum. In other words, when we go from a flat line to a curved circle, the energy levels are no longer evenly spaced. Therefore, our quantum oscillator exhibits an effective piston-like behavior through variations in the curvature parameter $\lambda$. In Fig. 2, we have plotted the dimensionless variation of energy gaps:
@@
as a function of the spatial curvature parameter. It is seen that for a fixed value of $\lambda$ $(\lambda\neq 0)$, the energy gaps increase with increasing the value of $n$. Furthermore, with increasing the curvature $\lambda$, the uniformity of the energy gaps between the eigen-energies diminishes further.
3 QUANTUM CURVATURE-DEPENDENT OTTO ENGINE
In this section, we consider a quantum Otto’s engine with a harmonic oscillator on a circle as its working substance, as illustrated in Fig. 3. This engine consists of two isochoric branches, one with a hot bath and the other with a cold thermal bath at two different temperatures $T_{h}>T_{c}$ [53, 54]. A quantum isochoric process (constant volume process) is similar to its classical isochoric process. In contrast, a classical adiabatic thermodynamic process does not always require the occupation probabilities to remain constant, while a quantum adiabatic process proceeds slowly enough so that the generic quantum adiabatic condition is satisfied, and keeps the population distributions constant [55]. Moreover, during quantum adiabatic stages, the working substance Hamiltonian varies while population distribution remains constant [56]. It is worth mentioning that if this variation in the Hamiltonian causes the energy gaps to scale at a constant ratio, the efficiency of the engine is like its classical counterpart. However, if the deformation of the Hamiltonian leads to inhomogeneous energy gaps, the heat engine’s efficiency can exceed its classical counterpart [57].
@@ 3 QUANTUM CURVATURE-DEPENDENT OTTO ENGINE
The first step in understanding the thermodynamics of a harmonic oscillator on a circle is to find the correct expression for the equilibrium distribution when the system is in contact with a thermal bath. The equilibrium distribution of a quantum oscillator on a circle, weakly coupled to a large classical bath, will eventually reach the corresponding Gibbs thermal equilibrium state, $\rho=e^{-\hat{H}(\lambda)/T}/Z(\lambda)$, [58] where the canonical partition function of the quantum oscillator, acting as the working substance of our Otto’s engine, is given by
@@
3.1 Processes of a quantum Otto engine cycle
@@ 3.1.1 Hot isochore $(A\rightarrow B)$
$Q_{h}=\sum_{n}E_{n}(\lambda)dP_{n}(\lambda)=\sum_{n}E_{n}^{A}(\lambda_{2})\Big{[}P_{n}^{B}(\lambda_{2})-P_{n}^{A}(\lambda_{2})\Big{]}.$ (8)
@@
3.1.2 Isentropic expansion $(B\rightarrow C)$
@@
$W_{B\rightarrow C}=\sum_{n}P_{n}dE_{n}=\sum_{n}P_{n}^{B}(\lambda_{2})\Big{[}E_{n}^{A}(\lambda_{2})-E_{n}^{C}(\lambda_{1})\Big{]}.$ (9)
3.1.3 Cold isochore $(C\rightarrow D)$
@@
$Q_{c}=\sum_{n}E_{n}^{C}(\lambda_{1})\Big{[}P_{n}^{D}(\lambda_{1})-P_{n}^{C}(\lambda_{1})\Big{]},$ (10)
@@
3.1.4 Isentropic compression $(D\rightarrow A)$
@@
$W_{D\rightarrow A}=\sum_{n}P_{n}^{D}(\lambda_{2})\Big{[}E_{n}^{D}(\lambda_{2})-E_{n}^{A}(\lambda_{1})\Big{]}.$ (11)
@@
3.2 Work and heat of the quantum Otto cycle
Now we can obtain the net work performed during the cycle [58] as follows:
$W$ $=$ $-(W_{B\rightarrow C}+W_{D\rightarrow A})=Q_{h}-Q_{c}$
$=$ $\sum_{n}\Big{[}E_{n}^{A}(\lambda_{2})-E_{n}^{C}(\lambda_{1})\Big{]}\Big{[}P_{n}^{B}(\lambda_{2})-P_{n}^{A}(\lambda_{2})\Big{]}.$
@@ 3.2 Work and heat of the quantum Otto cycle
4 THE THERMAL EFFICIENCY OF QUANTUM OTTO ENGINE
The first law of thermodynamics explains the conversion of heat, a form of energy, into work, while the second law determines how much of this heat can be effectively converted into work. Efficiency is defined as the ratio of output energy to input energy [60]. The thermal efficiency of our Otto cycle is defined by
@@ 4 THE THERMAL EFFICIENCY OF QUANTUM OTTO ENGINE
4.1 LIMITING CASE OF THERMAL EFFICIENCY
@@
4.1.1 The small spatial curvature limit
@@
4.1.2 The large spatial curvature limit
@@
$Q_{h}=\frac{\lambda}{2}\Big{[}\frac{q_{h}\theta^{\prime}_{3}(q_{h})}{\theta_{3}(q_{h})-1}-\frac{q_{c}\theta^{\prime}_{3}(q_{c})}{\theta_{3}(q_{c})-1}\Big{]},$
@@
5 RESULTS AND REMARKS
In this section, we proceed to study the effects of curvature on our quantum Otto’s engine using the analog model of a simple quantum harmonic oscillator on a circle. As shown in Eq. (3.2), the work extracted and thus the heat engine’s efficiency of our proposed quantum Otto’s heat engine depend on the net variation of the energy eigenvalues between points $A$ and $C$, as well as the population difference between points $A$ and $B$.
@@ 5 RESULTS AND REMARKS
One of the important factors in enhancing the efficiency of a quantum heat engine over it’s classical counterpart is that the change in the Hamiltonian of it’s working substance should be such that the difference between the energy gaps during the engine’s irreversible process changes in an anharmonicity manner [57].
It is worth noting that the results of Fig. 9 indicate that the efficiency of our analog heat engine model depends on the spatial curvature of the working substance’s location. Therefore, by utilizing a quantum harmonic oscillator on a circle as the working substance, we can control the work and efficiency of the Otto engine. Moreover, it is possible to enhance it’s efficiency to the Carnot bound $(\eta_{Carnot}=1-{T_{cold}}/{T_{hot}})$, by adjusting the curvature of space parameters $\lambda_{1}$ and $\lambda_{2}$ (the height of the hot and cold baths). It is worth mentioning that despite the increase in efficiency, our quantum Otto engine remains bound by the Carnot limit and fully adheres to the second law of thermodynamics [45].
Finally, to better visualize the results that have been obtained, we display in Fig. 10 the variation of our engine efficiency versus the curvatures $\lambda_{1}$ and $\lambda_{2}$.
@@
Figure 11: The heats $Q_{h}$ and $Q_{c}$ interchanged between the system and the thermal baths versus $\lambda_{2}$ for $\lambda_{1}=0.1$.
Remark: In our curvature-dependent quantum Otto engine, once the Carnot efficiency is achieved, further increasing the curvature difference between the two thermal baths leads to a transition from heat engine operation to refrigeration. For instance, Fig. 11 illustrates the heat exchanged between the working substance and both the cold and hot baths as a function of $\lambda_{2}$, with $\lambda_{1}=0.1$. As shown, when $\lambda_{2}$ exceeds approximately $7.2$, the heat exchanged with the hot bath becomes negative, while the heat exchanged with the cold bath remains positive, signifying the transition from a heat engine to a refrigerator.
6 SUMMARY AND CONCLUDING REMARKS
In this paper, we have considered a quantum Otto’s engine with a quantum harmonic oscillator on a circle as it’s working substance, serving as an analog model of general relativity. By utilizing the curvature-dependent energies of this oscillator, we have investigated how the curvature effects of the physical space affects the properties of our quantum heat engine. Assuming that two thermal baths are located at places with different curvatures, we have calculated the curvature-dependent work and heat, with a particular focus on the effects of curvature on it’s thermal efficiency. The result show that by adjusting the curvature difference between the bath’s locations, we can control the efficiency of our heat engine, potentially allowing it to reach the Carnot efficiency limit.
[…]
Acknowledgements